# Generalized Frobenius Manifold Structures on the Orbit Spaces of Affine Weyl Groups I

### v1

@@ 5. Examples
$D=\left(t_{1}^{4}-6t_{2}t_{1}^{2}+6t_{3}t_{1}+3t_{2}^{2}\right)\left(t_{1}^{4}+6t_{2}\ t_{1}^{2}+6t_{3}t_{1}+3t_{2}^{2}\right).$
We also have $e^{\flat}=-(\mathrm{d}\log z_{1}+\mathrm{d}\log z_{3})$. The monodromy group is given by
$\mathrm{Mono}(\mathcal{M}(A_{3},\omega_{1}+\omega_{3}))=\mathrm{Stab}(\omega_{1}+\omega_{3})=\langle\sigma_{2}\rangle\rtimes\mathbb{Z}^{3}\cong\mathbb{Z}_{2}\rtimes\mathbb{Z}^{2}.$
@@ 6. Conclusions
Given a root system $R$, a weight $\omega$ and a set of pencil generators $z^{1},\cdots,z^{\ell}$, we constructed a generalized Frobenius manifold structure $\mathcal{M}(R,\omega)$ on the orbit spaces of the associated affine Weyl groups $W_{a}(R)$. We proved that $z^{1},\cdots,z^{\ell}$ depend polynomially on the flat coordinates $t^{1},\cdots,t^{\ell}$ of the generalized Frobenius manifold, we show that the monodromy group $\mathrm{Mono}(\mathcal{M}(R,\omega))$ of the generalized Frobenius manifold can be embedded into the subgroup $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$ of $W_{a}(R)$.
In the subsequent paper [20], we will explicitly construct the pencil generators for affine Weyl groups of types $A_{\ell}$, $B_{\ell}$, $C_{\ell}$ and $D_{\ell}$ together with an appropriately chosen weight $\omega$, and compute the corresponding generalized Frobenius manifold structures in detail. For the case $(A_{\ell},\omega_{\ell})$, the generalized Frobenius manifold structure that we will present in [20] coincides with the one constructed in [7], where this generalized Frobenius manifold structure is constructed in terms of a super-potential given by a Fourier polynomial. It is shown in [7] that the bihamiltonian structure that is given by the associated flat pencil of metrics coincides with the dispersionless limit of the bihamiltonian structure of the q-deformed Gelfand-Dickey hierarchy [16, 17].
@@
    General construction of pencil generators: Can pencil generators be constructed for every affine Weyl group $W_{a}(R)$ and an arbitrary weight $\omega$?
@@
    Precise description of monodromy groups: While we have shown that $\mathrm{Mono}(\mathcal{M}(R,\omega))$ embeds into $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$, can this subgroup be characterized more explicitly?
@@
    Relation to other Frobenius manifold constructions: How do these generalized Frobenius manifolds relate to earlier constructions on the orbit spaces of the finite and the extended affine Weyl groups?
@@
    Almost duality and geometric interpretations: What is the Dubrovin almost-duality structure [11] for $\mathcal{M}(R,\omega)$?
@@
  6. vi)
    Associated integrable systems: Which integrable hierarchies already known in the theory of solitons are given by the topological deformations of the Principal Hiearchies [21, 22] of these generalized Frobenius manifolds?

### v3

@@ 5. Examples
$D=\left(t_{1}^{4}-6t_{1}^{2}t_{2}+6t_{1}t_{3}+3t_{2}^{2}\right)\left(t_{1}^{4}+6t_{1}^{2}t_{2}+6t_{1}t_{3}+3t_{2}^{2}\right).$
We also have $e=-\mathrm{grad}_{\eta}(\log z_{1}+\log z_{3})$.
Example 5.7 ($(R,\omega)=(D_{4},\omega_{2})$).
Let $R$ be the root system of type $D_{4}$, and $V=\mathbb{R}^{4}$. We take the simple roots
$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3},\quad\alpha_{3}=e_{3}-e_{4},\quad\alpha_{4}=e_{3}+e_{4}.$
Then we have the coroots $\alpha_{i}^{\vee}=\alpha_{i},~i=1,2,3$, and fundamental weights
$\omega_{1}=$ $~\alpha_{1}^{\vee}+\alpha_{2}^{\vee}+\frac{1}{2}\alpha_{3}^{\vee}+\frac{1}{2}\alpha_{4}^{\vee},\quad\omega_{2}=\alpha_{1}^{\vee}+2\alpha_{2}^{\vee}+\alpha_{3}^{\vee}+\alpha_{4}^{\vee},$
$\omega_{3}=$ $~\frac{1}{2}\alpha_{1}^{\vee}+\alpha_{2}^{\vee}+\alpha_{3}^{\vee}+\frac{1}{2}\alpha_{4}^{\vee},\quad\omega_{4}=\frac{1}{2}\alpha_{1}^{\vee}+\alpha_{2}^{\vee}+\frac{1}{2}\alpha_{3}^{\vee}+\alpha_{4}^{\vee}.$
Let us take $\omega=\omega_{2}$ and introduce coordinates $x_{1},x_{2}$ of $V$ by
$\mathbf{x}=c\,\omega_{2}+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2}+x_{3}\alpha^{\vee}_{3}+x_{4}\alpha^{\vee}_{4}.$
Since $\kappa=1$, we have $\lambda=\mathrm{e}^{-2\pi ic}$. We have the invariant $\lambda$-Fourier polynomials
$y_{1}=$ $~\lambda(\xi_{1}+\xi_{2}+\xi_{3}+\xi_{4}),$
$y_{2}=$ $~\lambda^{2}(\xi_{1}\xi_{2}+\xi_{1}\xi_{3}+\xi_{1}\xi_{4}+\xi_{2}\xi_{3}+\xi_{2}\xi_{4}+\xi_{3}\xi_{4}),$
$y_{3}=$ $~\frac{\lambda}{2}(\psi_{1}\psi_{2}\psi_{3}\psi_{4}-\phi_{1}\phi_{2}\phi_{3}\phi_{4}),$
$y_{4}=$ $~\frac{\lambda}{2}(\psi_{1}\psi_{2}\psi_{3}\psi_{4}+\phi_{1}\phi_{2}\phi_{3}\phi_{4}).$
Here we denote $v_{1}=x_{1}+c$, $v_{2}=x_{2}-x_{1}+c$, $v_{3}=x_{4}+x_{3}-x_{2}$, $v_{4}=x_{4}-x_{3}$, and
$\xi_{k}=\mathrm{e}^{2\pi iv_{k}}+\mathrm{e}^{-2\pi iv_{k}},\quad\psi_{k}=\mathrm{e}^{\pi iv_{k}}+\mathrm{e}^{-\pi iv_{k}},\quad\phi_{k}=\mathrm{e}^{\pi iv_{k}}-\mathrm{e}^{-\pi iv_{k}},\quad k=1,2,3,4.$
The metric on $V$ is given by
$\left((\alpha_{i}^{\vee},\alpha_{j}^{\vee})\right)=\left(\begin{array}[]{cccc}2&-1&0&0\\ -1&2&-1&-1\\ 0&-1&2&0\\ 0&-1&0&2\\ \end{array}\right),\quad\left(\mathrm{d}x_{i},\mathrm{d}x_{j}\right)=\left(\begin{array}[]{cccc}1&1&\frac{1}{2}&\frac{1}{2}\\ 1&2&1&1\\ \frac{1}{2}&1&1&\frac{1}{2}\\ \frac{1}{2}&1&\frac{1}{2}&1\\ \end{array}\right).$
In the coordinates $y_{1},\dots,y_{4}$ we have
$\left(g_{\lambda}^{ij}\right)=\frac{1}{4\pi^{2}}\left(\begin{array}[]{cccc}16\lambda^{2}-y_{1}^{2}+2y_{2}&3\lambda y_{3}y_{4}-y_{1}y_{2}&4\lambda y_{4}-\frac{1}{2}y_{1}y_{3}&4\lambda y_{3}-\frac{1}{2}y_{1}y_{4}\\ 3\lambda y_{3}y_{4}-y_{1}y_{2}&g_{\lambda}^{22}&3\lambda y_{1}y_{4}-y_{2}y_{3}&3\lambda y_{1}y_{3}-y_{2}y_{4}\\ 4\lambda y_{4}-\frac{1}{2}y_{1}y_{3}&3\lambda y_{1}y_{4}-y_{2}y_{3}&16\lambda^{2}-y_{3}^{2}+2y_{2}&4\lambda y_{1}-\frac{1}{2}y_{3}y_{4}\\ 4\lambda y_{3}-\frac{1}{2}y_{1}y_{4}&3\lambda y_{1}y_{3}-y_{2}y_{4}&4\lambda y_{1}-\frac{1}{2}y_{3}y_{4}&16\lambda^{2}-y_{4}^{2}+2y_{2}\\ \end{array}\right),$
where $g_{\lambda}^{22}=-64\lambda^{4}+4\lambda^{2}y_{1}^{2}+4\lambda^{2}y_{3}^{2}+4\lambda^{2}y_{4}^{2}-24\lambda^{2}y_{2}+2\lambda y_{1}y_{3}y_{4}-2y_{2}^{2}$. We introduce the pencil generators
$z_{1}=y_{1}+8\lambda,\quad z_{2}=y_{2}+24\lambda^{2}+\lambda(2y_{1}+2y_{3}-2y_{4}),\quad z_{3}=y_{3}+8\lambda,\quad z_{4}=y_{4}-8\lambda.$
In the new coordinates, we obtain a flat pencil of metrics $(g^{ij})+\lambda(\eta^{ij})$. The flat coordinates $t_{\alpha}$ of $\eta^{ij}$ are given by polynomials
$z_{1}$ $=\frac{8}{3}t_{1}^{4}+2t_{1}t_{4}+\frac{8}{3}t_{2}^{4}+2t_{2}t_{3},$
$z_{2}$ $=\frac{4}{9}t_{1}^{8}-\frac{4}{3}t_{1}^{5}t_{4}+\frac{56}{9}t_{1}^{4}t_{2}^{4}+\frac{20}{3}t_{1}^{4}t_{2}t_{3}+t_{1}^{2}t_{4}^{2}+\frac{20}{3}t_{1}t_{2}^{4}t_{4}+2t_{1}t_{2}t_{3}t_{4}+\frac{4}{9}t_{2}^{8}+t_{2}^{2}t_{3}^{2}-\frac{4}{3}t_{2}^{5}t_{3},$
$z_{3}$ $=-\frac{4}{3}t_{1}^{4}+8t_{1}^{2}t_{2}^{2}+2t_{1}t_{4}+2t_{2}t_{3}-\frac{4}{3}t_{2}^{4},$
$z_{4}$ $=\frac{4}{3}t_{1}^{4}+8t_{1}^{2}t_{2}^{2}-2t_{1}t_{4}+\frac{4}{3}t_{2}^{4}-2t_{2}t_{3}.$
In these coordinates the metric $\eta$ has the form
$\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}0&0&0&1\\ 0&0&1&0\\ 0&1&0&0\\ 1&0&0&0\end{pmatrix}$.
$F$ $=\frac{1}{t_{1}^{4}-t_{2}^{4}}\Bigg(-\frac{t_{1}^{12}}{1134}-\frac{t_{1}^{9}t_{4}}{180}+\frac{t_{1}^{8}t_{2}t_{3}}{36}-\frac{13t_{1}^{8}t_{2}^{4}}{1134}-\frac{t_{1}^{6}t_{4}^{2}}{24}+\frac{t_{1}^{5}t_{2}^{4}t_{4}}{30}-\frac{t_{1}^{5}t_{2}t_{3}t_{4}}{12}+\frac{t_{1}^{4}t_{3}^{3}}{96t_{2}}-\frac{t_{1}^{4}t_{2}^{2}t_{3}^{2}}{24}$
$\quad-\frac{t_{1}^{4}t_{2}^{5}t_{3}}{30}+\frac{13t_{1}^{4}t_{2}^{8}}{1134}-\frac{t_{1}^{3}t_{4}^{3}}{96}+\frac{t_{1}^{2}t_{2}t_{3}t_{4}^{2}}{16}+\frac{t_{1}^{2}t_{2}^{4}t_{4}^{2}}{24}+\frac{t_{1}t_{2}^{5}t_{3}t_{4}}{12}-\frac{t_{1}t_{2}^{2}t_{3}^{2}t_{4}}{16}-\frac{t_{1}t_{2}^{8}t_{4}}{36}$
$\quad+\frac{t_{2}^{3}t_{3}^{3}}{96}+\frac{t_{2}^{6}t_{3}^{2}}{24}+\frac{t_{2}^{9}t_{3}}{180}+\frac{t_{2}^{12}}{1134}-\frac{t_{2}^{4}t_{4}^{3}}{96t_{1}}\Bigg).$
the Euler vector field
$E=\frac{1}{4}t_{1}\frac{\partial}{\partial t_{1}}+\frac{1}{4}t_{2}\frac{\partial}{\partial t_{2}}+\frac{3}{4}t_{3}\frac{\partial}{\partial t_{3}}+\frac{3}{4}t_{4}\frac{\partial}{\partial t_{4}},$
and the unit vector field
$e=\frac{1}{D}$ $\left(6t_{1}\left(2t_{1}^{4}-3t_{4}t_{1}-10t_{2}^{4}-3t_{2}t_{3}\right)\frac{\partial}{\partial t_{1}}+6t_{2}\left(-10t_{1}^{4}-3t_{4}t_{1}+2t_{2}^{4}-3t_{2}t_{3}\right)\frac{\partial}{\partial t_{2}}\right$.
$-2\left(16t_{2}^{7}-30t_{3}t_{2}^{4}+112t_{1}^{4}t_{2}^{3}+120t_{1}t_{4}t_{2}^{3}+9t_{3}^{2}t_{2}+30t_{1}^{4}t_{3}+9t_{1}t_{3}t_{4}\right)\frac{\partial}{\partial t_{3}}$
$-2\left.\left(16t_{1}^{7}-30t_{4}t_{1}^{4}+112t_{2}^{4}t_{1}^{3}+120t_{2}t_{3}t_{1}^{3}+9t_{4}^{2}t_{1}+30t_{2}^{4}t_{4}+9t_{2}t_{3}t_{4}\right)\frac{\partial}{\partial t_{4}}\right),$
where
$D=4t_{1}^{8}+56t_{1}^{4}t_{2}^{4}+4t_{2}^{8}+60t_{1}^{4}t_{2}t_{3}-12t_{2}^{5}t_{3}+9t_{2}^{2}t_{3}^{2}-12t_{1}^{5}t_{4}+60t_{1}t_{2}^{4}t_{4}+18t_{1}t_{2}t_{3}t_{4}+9t_{1}^{2}t_{4}^{2}.$
We also have $e=-\mathrm{grad}_{\eta}\log z_{2}$.
@@ 6. Conclusions
Given a root system $R$, a weight $\omega$ and a set $\{z^{1},\dots,z^{\ell}\}$ of pencil generators, we present an approach to construct generalized Frobenius manifold structure $\mathcal{M}(R,\omega)$ on the orbit spaces of the associated affine Weyl groups $W_{a}(R)$, and we show that the monodromy group $\mathrm{Mono}(\mathcal{M}(R,\omega))$ of the generalized Frobenius manifold is isomorphic to the subgroup $\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell}$ of $W_{a}(R)$.
In the subsequent paper [20], we will explicitly construct the pencil generators for affine Weyl groups of types $A_{\ell}$, $B_{\ell}$, $C_{\ell}$ and $D_{\ell}$ together with appropriately chosen weights $\omega$, and compute the corresponding generalized Frobenius manifold structures in detail. For the case $(A_{\ell},\omega_{\ell})$, the generalized Frobenius manifold structure that we will present in [20] coincides with the one constructed in [7], where this generalized Frobenius manifold structure is constructed in terms of a super-potential given by a Fourier polynomial. It is shown in [7] that the bihamiltonian structure that is given by the associated flat pencil of metrics coincides with the dispersionless limit of the bihamiltonian structure of the q-deformed Gelfand-Dickey hierarchy [16, 17].
@@
    General construction of pencil generators: Can pencil generators be constructed for every affine Weyl group $W_{a}(R)$ and an arbitrarily chosen weight $\omega$?
@@
    Relation to other Frobenius manifold constructions: Are these generalized Frobenius manifolds related to the Frobenius manifolds constructed in [9, 12] on the orbit spaces of the Coxeter groups and the extended affine Weyl groups?
@@
    Almost duality: What is the Dubrovin almost-duality structure [11] for $\mathcal{M}(R,\omega)$?
@@
    Associated integrable systems: Which integrable hierarchies already known in the theory of integrable systems are given by the topological deformations of the Principal Hiearchies [21, 22] of these generalized Frobenius manifolds?
Appendix A Proof of Lemma 3.5
We are to prove that
$\mathrm{Re}(d_{\alpha})>0,\quad\alpha=1,\dots,\ell.$ (A.1)
To this end, we first show a general version of Theorem 4.1, the conclusion of which is not based on the assertion of Lemma 3.5. Since we do not assume a priori that all degrees $d_{\alpha}$ have positive real part, we may come across resonance cases when solving the differential equation of quasi-homogeneous periods, i.e., it could happen that
$d_{\beta}-d_{\alpha}=j\kappa,\quad j\in\mathbb{Z}$
for some $\alpha$ and $\beta$, so in such a case the quasi-homogeneous periods may contain logarithmic terms. In what follows we follow the notations of Sect. 4.
Lemma A.1.
For each quasi-homogeneous flat coordinate $t^{\alpha}=t^{\alpha}(z)$ of the metric $\eta$ with degree $d_{\alpha}$, there is a quasi-homogeneous period $u\in\mathcal{S}_{\Lambda}$ such that
$u=v^{\alpha}(z,\rho)=\rho^{d_{\alpha}/\kappa}h^{\alpha}(z,\rho),$ (A.2)
where $h^{\alpha}(z,\rho)$ has the form
$h^{\alpha}(z,\rho)=t^{\alpha}(z)+\sum_{j=1}^{\infty}\rho^{j}\Big(h^{\alpha}_{j,0}(z)+\log\rho~h^{\alpha}_{j,1}(z)\Big).$ (A.3)
Proof.
To find a solution $u$ of the form (A.2) for the equations (4.5) and (4.8), we need to solve the following system of equations:
$\frac{\partial\vec{\xi}}{\partial z^{i}}$ $=\Gamma_{i}(\rho)\vec{\xi},\quad i=1,\cdots,\ell,$ (A.4)
$\kappa\rho\frac{\partial\vec{\xi}}{\partial\rho}$ $=(A(\rho)-d_{\alpha})\vec{\xi},$ (A.5)
where
$\vec{\xi}=\left(\frac{\partial h^{\alpha}}{\partial z^{1}},\dots,\frac{\partial h^{\alpha}}{\partial z^{\ell}}\right)^{T},\quad(\Gamma_{i}(\rho))_{j}^{k}=\Gamma_{ij}^{k}(\rho).$
We expand $A(\rho)$ into a power series
$A(\rho)=A+\rho A_{1}+\rho^{2}A_{2}+\dots$
and expand $\vec{\xi}$ into the form given by (A.3) as
$\vec{\xi}=\vec{\xi}_{(0)}+\sum_{j=1}^{\infty}\rho^{j}\Big(\vec{\xi}_{(j,0)}+\log\rho~\vec{\xi}_{(j,1)}\Big).$
Then the equation (A.5) is equivalent to
$(A-d_{\alpha})\vec{\xi}_{(0)}$ $=0,$ (A.6)
$(A-d_{\alpha}-j\kappa)\vec{\xi}_{(j,1)}$ $=-\sum_{r=1}^{j-1}A_{r}\vec{\xi}_{(j-r,1)},\quad j=1,2,\dots,$ (A.7)
$(A-d_{\alpha}-j\kappa)\vec{\xi}_{(j,0)}$ $=\kappa\vec{\xi}_{(j,1)}-A_{j}\vec{\xi}_{(0)}-\sum_{r=1}^{j-1}A_{r}\vec{\xi}_{(j-r,0)},\quad j=1,2,\dots,$ (A.8)
here (A.6) is automatically satisfied. Note that $A$ is diagonalizable due to the assumption that $\nabla E$ is diagonalizable, hence for each $\zeta\in\mathbb{C}$, $\ker(A-\zeta)\oplus\mathrm{Im}(A-\zeta)=\mathbb{C}^{\ell}$. Therefore, for each $j=1,2,\dots$, whether $d_{\alpha}+j\kappa$ is an eigenvalue of $A$ or not, we can solve $\vec{\xi}_{(j,0)}$ and $\vec{\xi}_{(j,1)}$ uniquely from the equations (A.7) and (A.8). Then we obtain a solution $\vec{\xi}$ of (A.5) by induction on $j$. Since $\vec{\xi}_{(0)}$ solves (A.4), and equations (A.4) are compatible with (A.5), we know that $\vec{\xi}$ also solve (A.4). The lemma is proved. ∎
It follows from the above Lemma that for each index $\alpha$, there exist constants $C^{\alpha}_{\beta}$ and $S^{\alpha}$, such that
$v^{\alpha}(z,\rho)=C^{\alpha}_{\beta}\check{x}^{\beta}(z,\rho)+S^{\alpha}.$ (A.9)
Now we focus on the behaviour of the right hand side of (A.9) when $\rho\to 0$. Note that the Laurant polynomials $Z_{1},\cdots,Z_{\ell}$ given in (4.6) induce the canonical projection
$Z=(Z_{1},\cdots,Z_{\ell})\colon\mathbb{C}^{\ell}\to\mathbb{C}^{\ell}/W_{a}(R),$
where $\check{x}^{1},\dots,\check{x}^{\ell}$ are coordinates on $\mathbb{C}^{\ell}$ and the affine Weyl group $W_{a}(R)$ acts canonically on $\mathbb{C}^{\ell}=\mathbb{R}^{\ell}\otimes\mathbb{C}$. Denote the inverse map of $Z$ by $\check{X}$, then we have
$\check{x}^{\beta}(z,\rho)=\check{X}^{\beta}(\rho^{\theta_{1}/\kappa}z_{1},\cdots,\rho^{\theta_{\ell}/\kappa}z_{\ell}).$
Note that $Z$ is a branched covering map, hence there exists a neighbourhood $U$ of $(Z_{1},\dots,Z_{\ell})=(0,\dots,0)$ such that its preimage is a disjoint union of some open subsets of $\mathbb{C}^{\ell}$, on each one of which merely one point is mapped to $(Z_{1},\dots,Z_{\ell})=(0,\dots,0)$. Since $(\rho^{\theta_{1}/\kappa}z_{1},\dots,\rho^{\theta_{\ell}/\kappa}z_{\ell})$ lies in $U$ for sufficiently small $\rho$, we know that on each monodromy branch, $\check{x}(z,\rho)=(\check{x}^{1}(z,\rho),\dots,\check{x}^{\ell}(z,\rho))$ converges to a constant point when $\rho\to 0$, regardless of the value of $z=(z_{1},\dots,z_{\ell})$. In particular, this constant point is one of the common roots of $Z_{r}$ for $r=1,\cdots,\ell$. Therefore, the right hand side of (A.9) converges to a constant $V^{\alpha}$ when $\rho\to 0$ on each monodromy branch. It follows from Lemma A.1 and (A.9) that
$\rho^{d_{\alpha}/\kappa}\left(t^{\alpha}(z)+\sum_{j=1}^{\infty}\rho^{j}\Big(h^{\alpha}_{j,0}(z)+\log\rho~h^{\alpha}_{j,1}(z)\Big)\right)\to V^{\alpha},\quad\textrm{when}\ \rho\to 0.$
Since $t^{\alpha}(z)$ is not a constant, we have $V^{\alpha}=0$ and $\rho^{d_{\alpha}/\kappa}\to 0$ when $\rho\to 0$ on each monodromy branch. This implies that $\mathrm{Re}(d_{\alpha}/\kappa)>0$, hence $\mathrm{Re}(d_{\alpha})>0$. The theorem is proved.
Acknowledgement
This work is supported by NSFC No. 12571266.